My attempt to passively measure the transfer function of the foam failed fantastically.
As it turns out, the room temperature fluctuations inside the PSL box reach the 1 mK/rHz noise floor of the AD590 (or maybe the ADC) at ~1-2 mHz. Everything at higher frequencies is noise.
So to see what the foam is doing we will have to do something smarter - we need a volunteer to disable the RC temperature servo from the EPICS screen and then cycle the PSL table lights every hour in the morning.
We'll then use our knowledge of the Laplace transform to get the TF from the step responses.
more detailed instructions needed....